OpenStudy (anonymous):

If steel has a heat capacity of 0.452 kj/kg/C. And you need to go raise the temperature by 1350 C. How many joules of heat would you need?

OpenStudy (frostbite):

Hello @jellyfishdream This question is very simple, but in order to solve the problem we need a vital information, as you might have noticed your self: We need to know the mass of the steel sample we want to raise the temperature of.

OpenStudy (frostbite):

If we know the mass we can simply use the relation between the specific heat capacity: \[\Large E=c \times m \times \Delta T\]

OpenStudy (anonymous):

I know the Density is 7,874 Kg/m3

OpenStudy (anonymous):

which numbers get plugged into where?

OpenStudy (frostbite):

Well if we know the density, then we just get a new variable which is the volume, \(Q\) is the energy added as heat, \(c\) is the specific heat capacity, \(\Delta T\) is the difference in temperature. \(\rho\) is the density and \(V\) the volume

OpenStudy (frostbite):

Hmm then we just get (sorry I wrote wrong the first time) \[\Large Q=c \times \Delta T \times \rho \times V\]

OpenStudy (anonymous):

Q= 0.452 kj/kg/C x ΔT x 7,874 Kg/m3 x V

OpenStudy (anonymous):

Where do I plug in the 1350 C?

OpenStudy (frostbite):

in \(\Delta T\) as it is your increase in temperature.

OpenStudy (frostbite):

But you still need a volume then.

OpenStudy (frostbite):

As I said, we are lacking some information.

OpenStudy (frostbite):

Furthermore the density is usual temperature dependent. So you can also take in considerations if it is a fair approximation to assume the density remains constant when you increase the temperature with 1350 degrees C

OpenStudy (anonymous):

I was also given a melting point of: 1370 C but no volume

OpenStudy (frostbite):

Any chance you could take a pic of the question if it is from a book and attach it as a file? ect ect?

OpenStudy (frostbite):

It seems like something is completely wrong, or I am just not aware of a possible solution.

OpenStudy (anonymous):

I'm doing the extra credit. I don't really understand it though.

OpenStudy (frostbite):

Yeah okay.... hmmm well first of the ones who wrote these questions this are not super genius. They constantly write the "heat capacity" but they constantly write what is known as the "specific heat capacity" which we define as: \[\Large c _{s}=\frac{ C }{ m }\] \(c_{s}\) is the specific heat capacity, \(C\) is the heat capacity, \(m\) is the mass. From this definition, \(c_{s}\) gets the units joules per kelvin per gram (\(J K^{-1}g^{-1}\))

OpenStudy (frostbite):

So it seems to me you are lacking information to solve the problem and they haven't throughout the question.

OpenStudy (anonymous):

Yeah, sorry. My teacher kind of sucks at teaching chemistry obviously. I'm just giving up on the question.

OpenStudy (frostbite):

Don't be sorry @jellyfishdream I just want to give you some good arguments if you confront them with the assignment, then they might also get a chance to correct it.

OpenStudy (anonymous):

Thanks for everything @Frostbite